OpenStudy (anonymous):

Can someone check my answer please? Tell me which one of my answers are correct please?

OpenStudy (anonymous):

1. A container holds 15 pennies, 8 nickels, and 10 dimes. You will randomly select two coins without replacement. Show your work! a. How many ways can you select the coins? b. How many ways can you select exactly 1 nickel? c. What is the probability that you select 2 pennies? d. What is the probability that you select a dime and then a penny? Then, you should get... A. Fill in the probabilities on each branch of the tree diagram. (I don't get this part) https://www.flickr.com/photos/102326650 @N03/13744724835/ (Is this the right answer for a-d?) (15 + 8 + 10)! = 33! ways 1b. How many ways can you select exactly 1 nickel? Then, that would be 1 * (15 + 10)! = 25! ways 1c. What is the probability that you select 2 pennies? That is easy if you are only allowed to select up to 2 coins... P(2 pennies) = 15/33 * 14/32 1d. P(a dime and a penny) = 10/33 * 15/32 (Or is this the right answer for a-d?) Each fraction on the tree diagram will be found by multiplying the corresponding probabilities. For example, the answer at the top of the page is 15/33 * 14/32 = 35/176. a. 33C2 = 528 (not 33!) b. How many ways to select 1 nickel: (8C1)(25C1) = 8(25) = 200 c. Probability of selecting 2 pennies: (15C2)/(33C2) = 35/176 = .19886 d. Probability of dime then penny: 10/33 * 15/32 = 25/176 = .142

OpenStudy (campbell_st):

(a) well I would have thought 2 coins without replacement would have been 33 x 32 you choice any 1 of the 33 coins available... on the 1st selection and then anyone of the 32 coins remaining...

OpenStudy (campbell_st):

(b) is you select a nickel 1 nickel 8 x 8 you select any one of the 8 nickels on the 1st draw.... if you don't get a nickel on the 1st draw there are still 8 to choice from on the 2nd draw

OpenStudy (anonymous):

So neither of my answers are correct?

OpenStudy (campbell_st):

parts (c) and (d) are correct.... thats the way I see it... if you look at the probabilities for (c) and (d) you are saying the number of possible outcomes is 33 x 32 but you answer for (a) says you are picking 33! ways.... here is the tree diagram  hope it helps

OpenStudy (campbell_st):